OpenStudy (anonymous):

A projectile is thrown upward so that its distance is above the ground after t seconds is given by the function h(t)=-16t^2+576t. After how many seconds does the projectile reach its maximum height?

OpenStudy (anonymous):

Kirby can you help me? @kirbykirby

OpenStudy (kirbykirby):

oh yes sorry I got distracted by another question.. I'm almost done typing

OpenStudy (kirbykirby):

If you don't know calculus, then you can do this knowing that this is a quadratic function, so the maximum height occurs at the vertex (this is because you have a parabola of the forma \(at^2+bt+c\), where \(a<0\) (here c = 0), so the parabola looks like this \(\cap\). So an easy way to handle this is to manipulate your equation into vertex form: \(-16t^2+576 t \\ =-16(t^2-36t)\) complete the square by adding and subtracting \(\left(\frac{b}{2}\right)^2=\left(\frac{-36}{2}\right)^2=324\) \(=-16(t^2-36t+324-324)\\ =-16(t-18)^2-16(-324)\\ =-16(t-18)^2+5184\) Your parabola is of the form \(a(x-h)^2+k\) now, so the hack occurs at the vertex. So the maximum height is 5184 at t=18

OpenStudy (kirbykirby):

so the max occurs at the vertex*

OpenStudy (anonymous):

thank you so much, I actually found a older post of the same question where ganeshie8 answered exactly the same! Yours is clearer though and I really appreciate the help!

OpenStudy (kirbykirby):

oh awesome :) ! your welcome =]